OpenStudy (anonymous):

Prove that y = x/(x+c) is a general solution for the differntial equation dy/dx = (y - y^2)/x and show that all solutions contain (0,0).

OpenStudy (anonymous):

I've gotten to this so far, but I've no clue if it's right.

OpenStudy (anonymous):

\[ y(x)=\frac{x}{c+x}\\ y'(x)=\frac{c}{(c+x)^2}\\ \frac{y(x)-y(x)^2}{x}=\frac{\frac{x}{c+x}-\frac{x^2}{(c+x)^2}}{x }=\frac{c}{(c+x)^2}=y'(x) \]

OpenStudy (anonymous):

Thank you, I think I've got it now.

OpenStudy (anonymous):

You also can solve it directly. It is a separable differential equation.